OpenStudy (anonymous):

Mitchell gets an annual salary of $48 630.40.If she recives 17 1/2 % hloiday loading on the 4 week holiday period, calcaute: a)her normal pay for 4 weeks b) her holiday pay c) her holiday pay for 4 weeks please i need working out but simply

OpenStudy (anonymous):

o need help please

OpenStudy (anonymous):

to calculate her weekly pay divide the 48630.40 by 52. this is her weekly pay, then multiply that by 4 for her normal pay.

OpenStudy (anonymous):

her weekly pay is roughly 935.19, then multiply that by 4 to get her monthly pay (A),

OpenStudy (anonymous):

so what the equation please

OpenStudy (anonymous):

for a b c

OpenStudy (anonymous):

48630.40/52=1/4x thats A

OpenStudy (anonymous):

take what you got from A (before you multiplied by 4) and plug it into this .175x=B

OpenStudy (anonymous):

C=B+A

OpenStudy (anonymous):

i dont understand i need the equation with the answer please